The winning ticket of a lottery is a sequence of three numbers chosen from the numbers 1 through 130. (Enter your probabilities as fractions.)

(a) If a number may repeat in the sequence, find the probability that a player guesses the winning sequence.
(b) If a number may not repeat in the winning sequence, find the probability that a player guesses the winning sequence.

Answer by Boreal(15235)

the second is 3/130*2/129*1/128=6/2,146,560, which can be reduced to 1/357760.
the first is (1/130)^3=1/2,197,000